1) (4 pts) Black hair in rabbits is determined by a dominant allele, B, and white hair by homozygosity for a recessive allele, b. Two heterozygous rabbits mate and produce a litter of seven offspring. With what frequency would we expect to find:

  1. a) four whites and three black offspring?

  2. b) one white and six black offspring?
Answer #1

a).

Bb x Bb ------Parents

B

B

B

BB (black)

Bb (black)

b

Bb (black)

bb (white)

Black = ¾

White = 1/4

n!/s!t! (a)^s(b)^t

Expected proportion of black = a = ¾

Expected proportion of white = b = 1/4

Observed Black = s = 3

Observed white =t = 4

n = total = 7

P = (7! / (3! X 4!)) (3/4)^3 (1/4)^4

= (5040/3*24) (27/64)(1/256)

=70*0.42*0.004

= 0.11

Probability of 4 white and 3 black offspring = 0.11

b).

P = (7! / (6! X 1!)) (3/4)^6 (1/4)^1

= (5040/720) (0.178)(0.25)

=7*0.178*0.25

= 0.312

Probability of 1 white and 6 black offspring = 0.312